# Implemetation of a shooting technique for quantum optimal control on spin qudits

### 23 Mar 2026 v1

@@ II.2.1 Pontryagin’s Maximum Principle (PMP)
The MAGICARP is a PMP inspired method, a mathematical tool to reframe an optimisation problem on “trajectories” (living in an infinite dimensional space) into an Hamiltonian problem where only a finite amount of coefficients are to be optimized. This change from point of view is analogous to using a principle of least action in classical physics. The problem of optimisation of the control fields can be rewritten as a problem where
@@
with $u(t)$ a vector with all the $u_{k}(t)$ with $k\leq N_{\text{ctrls}}\in\mathbb{N}$, $U(t)$ the gate obtained at time t via the optimisation protocol, $U_{0}$ that is usually the identity and $U_{f}$ the target gate. We also have a cost to minimize during this procedure
$\mathcal{C}=G(U(t_{f}),t_{f})+\int_{0}^{t_{f}}F_{0}(U(t),u(t),t)\,dt,$ (14)
$G$ could be $t_{f}$ plus the fidelity times an arbitrary constant for instance. In $F_{0}$ we could include the fact that the pulse amplitude has to be maximized with $F_{0}=-\sqrt{\sum_{k}{u_{k}(t)}^{2}}$ for instance. If the pulse amplitude ends up being higher than $\Omega_{\text{Rabi}}$ the maximum amplitude, we could always renormalize the controls by applying them for a longer time like we did for GRAPE thanks to the equivalence between amplitude and time explained in section II.1.1 and detailed in section II.1.3.
The Pontryagin’s Maximum Principle [29] is an important tool of optimal control theory [1]. Starting from static initial and final (objective) quantum gates (That is static boundary conditions), it turns the optimization of our problem relatively to a cost function into the problem of minimizing an action that is the following:
@@
Let $U(t)\in\mathbb{R}^{n\times n}$ be the quantum gate we are constructing, at time $t$, $\Lambda(t)\in\mathbb{R}^{n\times n}$ an adjoint state, defining the optimisation we choose for the problem., and $u(t)\subset\mathbb{R}^{N_{\text{ctrls}}}$ the controls $u_{k}(t)$ with $k\leq N_{\text{ctrls}}\in\mathbb{N}$ in eq. 8. Let $G$ be a terminal cost function, for instance the fidelity, $F_{0}$ a running cost and $F$ a vector function describing the system dynamics. The optimal control is defined as the one minimizing the action $S$:
@@
The function under the integral is the Lagrangian $\mathcal{L}$ of the optimal control problem and $\Lambda(t)$ is an adjoint state. The optimal control is then defined as an extremum of the action $S$. We can write it as
@@
$+\left(\frac{\partial G}{\partial U(t_{f})}+\Lambda(t_{f})\right)\delta U(t_{f})+\Lambda(0)\delta U(0).$
By integrating by part the term $\Lambda\delta\dot{U}$. Thus the extremals of $S$ fulfill the following conditions
Property II.2.
The optimal solution of the PMP follows:
$\dot{\Lambda}$ $=\frac{\partial F_{0}}{\partial U}-\Lambda\frac{\partial F}{\partial U}$ (15)
$\dot{U}$ $=F$ (16)
$\frac{\partial F_{0}}{\partial u}$ $=\Lambda\frac{\partial F}{\partial u}.$ (17)
$U(0)=U_{0}$ $,\quad U(t_{f})=U_{f}$
From a physical point of view eq. 15 describes a Euler-Lagrange equation for $U$, eq. 16 describes the dynamics of the system (a Euler-Lagrange equation for $\Lambda$) and eq. 17 imposes a condition on the controls (a Euler-Lagrange equation for $u$). It is then possible to translate this formulation into an Hamiltonian problem, with the Pontryagin Hamiltonian defined as:
$H_{P}$ $=\Lambda\cdot\dot{U}-\mathcal{L}=\Lambda\cdot F-F_{0}$
With $\cdot$ denoting the Hilbert–Schmidt inner product. Given that we will follow a physcial trajectory of the system where eq. 16 is satisfied. The solution to the optimization problem is then written as
@@ II.2.2 The MAGICARP
When inputing the running cost $F_{0}=-\sqrt{\sum_{k}{u_{k}(t)}^{2}}$ in the Hamiltonian optimization problem of eq. 18 Jankovic derived in [15]
Property II.3.
Pulses $u_{k}(t)$ of eq. 8 with $k\leq N_{\text{ctrls}}\in\mathbb{N}$ that minimize the running cost $F_{0}$ are of the form
$\frac{u_{k}(t)}{c(t)}=\operatorname{Re}\operatorname{Tr}\left(U(t)MU^{\dagger}(t)H_{k}\right),$ (22)
where
$c(t)=\sqrt{\sum_{k}u_{k}^{2}(t)},$ (23)
@@
$\dot{\Lambda}$ $=iH\Lambda,$
$\dot{U}$ $=\frac{\partial H_{P}}{\partial\Lambda},$ (26)
@@
Then a very useful property to notice is that eq. 26 is easy to solve. Since $U\in SU(d)$ with $d$ the dimension of the Hilbert space (up to a global phase), then $\Lambda$ is an element of the tangent space of $SU(d)$ at $U$, given by the Lie algebra $\mathfrak{su}(d)$ multiplied by $U$.
$\Lambda(t)=U(t)\Lambda_{0}(t),$ (28)
and since $SU(d)$ is a Lie group $\Lambda_{0}(t)$ is constant in time. Then we can rewrite $i\Lambda^{\dagger}$ in eq. 27 as $MU^{\dagger}$ with $M$ a constant, traceless, Hermitian matrix. Finally by using the cyclic property of the trace and $\frac{\partial H}{\partial u_{k}}=H_{k}$ for all $k$ we obtain eq. 22. ∎
We could see the $c(t)$ as constant in time if we maximize the amplitude of the controls by this quantity when there is no drift. The MAGICARP is about optimizing the constant coefficients of the matrix $M$ to minimize the cost function in eq. 9. In the following we will use a gradient descent for this purpose.
@@ III.1.1 SMM
Here we present a bisphthalocyaninato terbium(III) ($\text{Tb}\text{Pc}_{2}$) molecule, also called a double decker, a SMM discovered in 2003 [14]. It is visible in the fig. 1a. The $\text{Tb}^{3+}$ ion has an effective nuclear spin of $I=3/2$ [11, 16], which gives rise to $2I+1=4$ energy levels. This makes the $\text{Tb}\text{Pc}_{2}$ an effective qu-4-it, a qudit [34] with $d=4$. It is also possible to superpose two double deckers into a triple decker $\text{Tb}_{2}\text{Pc}_{3}$ molecule, which then act as an effective qu-16-it [2] thanks to a coupling between the two $\text{Tb}^{3+}$ ions [19].
@@
Figure 1: (a) A double decker, $\text{Tb}\text{Pc}_{2}$. The blue atoms are some nitrogen and the grey ones are carbons. Image from [13]. (b) A triple decker, $\text{Tb}_{2}\text{Pc}_{3}$. The two dark blue atoms are still some $\text{Tb}^{3+}$ ions. Image from [13].
@@
To construct quantum gates we need to make transitions between energy levels of the system. Not all transitions are allowed. In the example of the double decker, selection rules only allows for adressing transitions between levels separated by a $I_{\pm}$ [16].
@@
Figure 2: (a) The energy levels graph of a double decker $\text{Tb}\text{Pc}_{2}$. The spacing is unequal thanks to a non linear term in the Hamiltonian. Otherwise energy levels would be positioned on the dashed lines. The resulting graph is a linearly coupled one. Values from [11]. (b) The energy levels graph of a triple decker $\text{Tb}_{2}\text{Pc}_{3}$. Each node represent an energy level, ranked by their energy. The edges represent adressable rabi transitions between nodes, for a certain energy difference in Kelvin. The energy levels are denoted by “lvl” with the first digit denoting the nuclear energy level for the first molecular magnet and the second digit the other nuclear energy level. This figure is done using parameters that differ slightly from [2], and at a high magnetic field to suppress level mixing.
@@
That means that only transitions shown in fig. 2a have an effect on the system. For instance, an energy level corresponding to I=1/2I=-1/2 can communicate when excited at the right frequency with the energy levels corresponding to I=3/2I=-3/2 or $I=1/2$ but not with the level corresponding to $I=3/2$. Thus only neighbouring interactions are allowed, resulting in what we call a linear graph. For the case of the triple decker, only coupling between levels separated by a $I_{1}$$_{\pm}$ or $I_{2}$$_{\pm}$ are allowed, the 1 or 2 indices denoting the nuclear energy levels for each of the $\text{Tb}^{3+}$ ions. The non trivial resulting graph is shown in fig. 2b. Usually the allowed transitions are determined experimentally and known to the experimentators. As long as all energy levels are connected, it is possible to create any unitary gate on the Hilbert space spanned by them.
@@ III.1.2 Running a Quantum Fourier Transform (QFT) on a SMM
The QFT is a key gate for quantum computation, including Shor’s algorithm and many other important algorithms, because it is often a good starting point for solving problems in the hidden subgroup problem [26, 5.4.3], a class of problems for which quantum computations can provide exponential speed ups. It is also referred in the literature as a generalized Hadamard gate on qudits.
@@
[…]
We show in fig. 4a the pulses used by the GRAPE method to get a QFT gate, and the population dynamics for this gate, starting from a maximally entangled state in fig. 4b.
[…]
We show in fig. 5a the pulses used by the MAGICARP to get a QFT gate, and the population dynamics for this gate, starting from a maximally entangled state in fig. 5b. One can notice that the amplitudes in fig. 5a are smoother than for fig. 4a. Although it is actually possible to take a large number of steps in the GRAPE method to have something smoother, it is still common for the GRAPE method to have frequent and sudden changes in the amplitudes (almost bang-bang like controls). This may be a problem for implementing those pulses experimentally. The MAGICARP avoids this kind of behavior on the examples we worked with, by deriving smooth controls.
@@ III.2 Numerical results
The MAGICARP allows for optimizing only a finite number of parameters (the constant coefficients of the matrix $M$ in eq. 22) rather than the infinite dimensional space of the control $u_{k}(t)$, that are smooth functions of time. An interesting property of this method is that the controls obtained fullfill a necessary (but no sufficient) condition for a time optimal control.
@@
Figure 6: Comparison of GRAPE and MAGICARP execution times. (a) Time to implement 1000 random Haar unitaries with precision $10^{-4}$ for a $\text{Tb}\text{Pc}_{2}$ molecule. (b) Time to implement 1000 random Haar unitaries with precision $10^{-4}$ for a $\text{Tb}_{2}\text{Pc}_{3}$ molecule.
@@
This restriction offers shorter gate execution times, especially for systems having a relatively large Hilbert space like the triple decker in fig. 6b. An important point is that it is increasingly computationally difficult to optimize the $M$ matrix of eq. 22 with the growing size of the Hilbert space. The natural gradient method developed in appendix C allows for finding solutions with sufficient fidelity (1e-4) at relatively high dimension (the triple decker example is of total Hilbert space dimension of 16). We used it to obtain the results of fig. 6b. Overall, the higher the dimension of the system is, the better the results of the MAGICARP in comparison to GRAPE when comparing median implementation times for random gates or the minimum implementation time for a specific gate (the QFT) as shown in fig. 7.
@@
Figure 7: Comparison of GRAPE and MAGICARP execution times on linear graphs with 3 to 9 nodes, as a function of the Hilbert space dimension. There is at least 100 initializations for each dimension. We used the natural gradient method described in appendix C, and an automatic differentiation with a L-BFGS-B gradient from SciPy [33] at low dimensions. The ratios indicate how much faster MAGICARP is compared to GRAPE for a given dimension. (a) Ratio of the median execution time for random Haar gates using GRAPE with random initializations over the median time using MAGICARP. (b) Ratio of the minimum execution time for a QFT gate using GRAPE with random initializations over the minimum time using MAGICARP.
@@
We choose to use a QFT gate to compare the minimum implementation for the MAGICARP and the GRAPE method since this gate is not trivial, has an implementation profile similar to a random gate as one can see when comparing fig. 6b and fig. 8a, and this gate is key to many important quantum algorithms. Note that for small dimensions (3 and 4), their is no improvement of the execution time with the natural gradient method used here since the ratios are under 1 in fig. 7 for the corresponding curve. Though if we compare the MAGICARP with the same gradient method as GRAPE (A L-BFGS-B gradient from SciPy [6, 33].), with the same parameters, and using automatic differentiation, we can compute the results at low dimensions (up to 5) and do witness that the MAGICARP is at least faster than GRAPE both in median time for random gates and in minimum time for the QFT gate for this naive implementation. It would be interesting to understand better how some gradient descents can ameliorate the MAGICARP.
Trying to do a specific gate with the best implementation times by trying many random initial conditions of our optimization algorithms is a method to get a better implementation time. We compared the MAGICARP with the GRAPE method for a QFT gate, T, X, and SUMX gate (a generalization of a CNOT on a qudit) as defined in [31] on a $\text{Tb}_{2}\text{Pc}_{3}$ molecule and obtained clearly better results in fig. 8, regarding both median and minimum time implementation.
@@
Figure 8: Comparison of GRAPE and MAGICARP execution times for 1000 random initializations with precision $10^{-4}$ on a $\text{Tb}_{2}\text{Pc}_{3}$ molecule. (a) QFT gate. (b) T gate. (c) X gate. (d) SUMX gate.

### 10 Apr 2026 v2

@@ II.2.1 Pontryagin’s Maximum Principle (PMP)
The MAGICARP consists in a finite-dimensional parametrization of controls in feedback form which is derived from the Pontryagin maximum principle (PMP), a mathematical tool to reframe an optimisation problem on “trajectories” (living in an infinite dimensional space) into an Hamiltonian problem where only a finite amount of coefficients, encoding the trajectories, are to be optimized. This change of point of view is analogous to using a principle of least action in classical physics. The problem of optimisation of the control fields can be rewritten as a problem where
@@
with $u(t)$ a vector of all the $u_{k}(t)$ for $k\leq N_{\text{ctrls}}\in\mathbb{N}$, $U(t)$ the gate obtained at time $t$ via the optimisation protocol, $U_{0}$ that is usually the identity and $U_{f}$ the target gate. To select a control among all those steering the system between the desired states, we seek to minimize a cost function $\mathcal{C}$ including a terminal cost $G$ and a running cost $F_{0}$ reflecting what a desirable control is:
$\mathcal{C}=G(U(t_{f}),t_{f})+\int_{0}^{t_{f}}F_{0}(U(t),u(t),t)\,dt.$ (30)
To achieve the transfer as quickly as possible, $G$ could either be the duration $t_{f}$ or, if we drop the hard constraint $U(t_{f})=U_{f}$, the infidelity between $U(t_{f})$ and $U_{f}$. In $F_{0}$ we could include the fact that the pulse amplitude has to be maximized with $F_{0}=-\sqrt{\sum_{k}{u_{k}(t)}^{2}}$ for instance. If the pulse amplitude ends up being higher than $\Omega$ the maximum amplitude of the pulses, we could always renormalize the controls by applying them for a longer time like we did for GRAPE thanks to the equivalence between amplitude and time explained in section II.1.1 and detailed in section II.1.3.
The PMP [32] is an important tool of optimal control theory. To strengthen the physical intuition behind it, we provide a heuristic derivation inspired from [3] of the corresponding necessary conditions of optimality by connecting it to the Euler-Lagrange equations from the calculus of variations. Starting from static initial and final target quantum gates (That is, static boundary conditions), it turns the problem of optimizing a cost function into the problem of minimizing an action
@@
Let $U(t)\in\mathbb{R}^{n\times n}$ be the quantum gate we are constructing, at time $t$, $\Lambda(t)\in\mathbb{R}^{n\times n}$ an adjoint state, defining the optimisation we choose for the problem., and $u(t)\subset\mathbb{R}^{N_{\rm ctrls}}$ the controls $u_{k}(t)$ with $k\leq N_{\rm ctrls}\in\mathbb{N}$ in eq. 24. Let $G$ be a terminal cost function, for instance the fidelity, $F_{0}$ a running cost and $F$ a vector function describing the system dynamics. The optimal control is defined as the one minimizing the action $S$:
@@
The function under the integral is the Lagrangian $\mathcal{L}$ of the optimal control problem and $\Lambda(t)$ is an adjoint state. The optimal control is then defined as a minimum of the action $S$. The first variation $\delta S$ of the action in response to a perturbation $\delta u$ of the control is formally given in terms of the state variation $\delta U$ by
@@
$+\left(\frac{\partial G}{\partial U(t_{f})}+\Lambda(t_{f})\right)\delta U(t_{f})+\Lambda(0)\delta U(0),$
by integration by parts of the term $\Lambda\delta\dot{U}$. Thus the minimums of $S$ fulfill the following conditions
Proposition II.2.
If a control $u$ associated to the state $U$ is optimal, then there exists an adjoint state $\Lambda$ such that the following conditions hold:
$\dot{\Lambda}$ $=\frac{\partial F_{0}}{\partial U}-\Lambda\frac{\partial F}{\partial U},$ (31)
$\dot{U}$ $=F,$ (32)
$\frac{\partial F_{0}}{\partial u}$ $=\Lambda\frac{\partial F}{\partial u},$ (33)
$U(0)=U_{0}$ $,\quad U(t_{f})=U_{f}.$
From a physical point of view, eq. 31 describes an Euler-Lagrange equation for $U$, eq. 32 describes the dynamics of the system (an Euler-Lagrange equation for $\Lambda$) and eq. 33 imposes a condition on the controls (a Euler-Lagrange equation for $u$). It is then possible to translate this formulation into an Hamiltonian problem, with the Pontryagin Hamiltonian defined as:
$H_{P}$ $=\Lambda\cdot\dot{U}-\mathcal{L}=\Lambda\cdot F-F_{0}.$
With $\cdot$ denoting the Hilbert–Schmidt inner product. Given that we will follow a physcial trajectory of the system where eq. 32 is satisfied, the solution to the optimization problem is then written as
@@ II.2.2 The MAGICARP
For a quantum evolution the eq. 32 corresponds to the Schrödinger equation and $U_{0}$ in eq. 38 is the identity. Inputing the running cost $F_{0}=-\sqrt{\sum_{k}{u_{k}^{2}(t)}}$ in the Hamiltonian optimization problem of eq. 34 Jankovic derived in [18]
Proposition II.3.
Pulses $u_{k}(t)$ of eq. 24 with $k\leq N_{\rm ctrls}\in\mathbb{N}$ that minimize the running cost $F_{0}$ are of the form
$u_{k}(t)=\operatorname{Re}\operatorname{Tr}\left(U(t)MU^{\dagger}(t)H_{k}\right),$ (39)
@@
$\dot{\Lambda}$ $=-iH\Lambda,$
$\dot{U}$ $=-iHU,$ (42)
@@
Then a very useful proposition to notice is that eq. 42 is easy to solve. By differentiating the function $\Lambda(t)^{\dagger}U(t)$, we find that its derivative is everywhere zero. Hence $\Lambda(t)=U(t)\Lambda(0)$. Since $\Lambda(0)$ is in the Lie algebra of the special unitary group, $\Lambda_{0}=iM$ for some traceless Hermitian matrix $M$. Then we can rewrite $i\Lambda^{\dagger}$ in eq. 43 as $MU^{\dagger}$ with $M$ a constant, traceless, Hermitian matrix. Finally by using the cyclic proposition of the trace and $\frac{\partial H}{\partial u_{k}}=H_{k}$ for all $k$ we obtain
$\frac{u_{k}(t)}{c(t)}=\operatorname{Re}\operatorname{Tr}\left(U(t)MU^{\dagger}(t)H_{k}\right),$
where
$c(t)=\sqrt{\sum_{k}u_{k}^{2}(t)},$
We could see the c(t) as constant in time if we maximize the amplitude of the controls by this quantity when there is no drift, so we end up with the expression of the optimal controls given in eq. 39. ∎
MAGICARP is a parametrization of controls by means of the initial matrix $M$. By means of a suitable choice of optimization routine, starting from an initial guess, this parameter is iteratively updated until the target gate is attained up to some user-defined tolerance. As such, it belongs to the class of shooting methods in which, roughly, one aims at finding the appropriate initial covector in such a way that the optimal set of equations derived from the PMP are satisfied while the desired target state is attained with sufficient precision. In particular MAGICARP suffers from the necessity of finding relevant initial guesses for the parameter $M$. However what is worth noting here is that, by working with the propagator or flow of the equation of motion and exploiting the Lie group structure, the maximization condition of the PMP leads to a control in closed-loop form. In other words, the control amplitudes are determined only by the instantaneous propagator $U(t)$. One notable advantage is that the parameter $M$ is not tied to any choice of discretization scheme and fully determines the associated control given the control Hamiltonians.
Remark 1.
Proposition II.3 is only concerned with the so-called normal extremals of the Pontryagin maximum principle. Put differently, for some choices of control Hamiltonians and some target gates, the optimal controls might not be given by Equation (39) for some traceless Hermitian matrix $M$. Nevertheless it appears natural to exploit numerically this simple parametrization in an attempt to obtain more time-efficient controls pulses. See Remark 2 in Appendix C for a discussion.
@@ III.1 The physics of SMM
The concept of Single-Molecule Magnets used as quantum devices started in the 90s [36, 23]. Their interest lies in their very small size and the fact that they show quantum effects. We will compare numerical results obtained via the MAGICARP on these two molecules.
@@ III.1.1 SMM
The first molecule of interest is the bisphthalocyaninato terbium(III) ($\text{Tb}\text{Pc}_{2}$) molecule, a SMM synthesised in 2003 [17] and shown in the fig. 1a which is also known as the double decker. The $\text{Tb}^{3+}$ ion has a nuclear spin of $I=3/2$ [14, 19], which gives rise to $2I+1=4$ energy levels. This makes the $\text{Tb}\text{Pc}_{2}$ an effective qu-4-it, a qudit [37] with $d=4$. It is also possible to superpose two double deckers into a triple decker $\text{Tb}_{2}\text{Pc}_{3}$ molecule, which then acts as an effective qu-16-it [4] thanks to a coupling between the two $\text{Tb}^{3+}$ ions [22].
@@
Figure 1: (a) A $\text{Tb}\text{Pc}_{2}$ double decker, composed of the central $\text{Tb}_{3+}$ ion (red) surrounded by nitrogen (blue) and carbon (grey) atoms of the phthalocynanine ligands. Image from [16]. (b) A triple decker, $\text{Tb}_{2}\text{Pc}_{3}$. The two dark blue atoms are also $\text{Tb}^{3+}$ ions. Image from [16].
@@
To construct quantum gates we need to induce transitions between the system’s energy levels by applying an electromagnetic pulse with an appropriate frequency. Not all transitions are allowed. In the example of the double decker, selection rules only allows for addressing transitions between levels separated by a $\Delta I=\pm 1$ [19].
@@
Figure 2: (a) The linear graph connecting the four energy levels of a double decker $\text{Tb}\text{Pc}_{2}$. The spacing is unequal thanks to a non linear term in the Hamiltonian. Otherwise energy levels would be positioned on the dashed lines. Values from [14]. (b) The graph connecting the sixteen energy levels of a triple decker $\text{Tb}_{2}\text{Pc}_{3}$. Each node represent an energy level, ranked by their energy. The edges represent addressable Rabi transitions between nodes, for a certain energy difference in Kelvin. The energy levels are denoted by “lvl” with the first digit denoting the nuclear energy level for the first molecular magnet and the second digit the other nuclear energy level. This figure is done using parameters that differ slightly from [4], and at a large magnetic field to suppress level mixing.
@@
This means that only transitions shown in fig. 2a have an effect on the system. For instance, an energy level corresponding to I=1/2I=-1/2 can communicate when excited at the right frequency with the energy levels corresponding to I=3/2I=-3/2 or $I=1/2$ but not with the level corresponding to $I=3/2$. Thus only neighbouring interactions are allowed, resulting in what we refer to as a linear graph. The resonances can be individually adressed thanks to a non-linear term that results from an electric quadrupolar interaction. For the case of the triple decker, only coupling between adjacent levels of nuclear spin $I_{1}$ and $I_{2}$ ($\Delta I=1$) are allowed, the 1 or 2 indices denoting the nuclear energy levels for each of the $\text{Tb}^{3+}$ ions. The non trivial resulting graph is shown in fig. 2b. Usually the allowed transitions are determined experimentally and as long as all energy levels are connected, it is possible to create any unitary gate on the Hilbert space associated to these SMM, for instance by using the GRD.
@@ III.1.2 Running a Quantum Fourier Transform (QFT) on a SMM
Figure 3: Control amplitudes (left panels) and population dynamics (right panels) for the implementation of a QFT gate on the double decker graph. Panels (a)–(b) correspond to GRD, (c)–(d) to GRAPE, and (e)–(f) to MAGICARP. In (a), each blue activation, delimited by black lines, correspond to a different $(\theta,\phi)$ for a GR in eq. 5. The length of each blue activation is $\phi$. For GRAPE and MAGICARP, each color denotes a distinct control amplitude. The target fidelity is $10^{-4}$. GRAPE controls are discretized in 50 time steps. All times are given in units of $\Omega^{-1}$ (see section II.1.3).
@@ III.2 Numerical results
For both the GRAPE and MAGICARP methods, we set the target infidelity to $10^{-4}$ in this work. We explain in section A.2 why this is a convenient choice. The MAGICARP allows optimization of only a finite number of parameters (the constant coefficients of the matrix $M$ in eq. 39) rather than the infinite dimensional space of the control $u_{k}(t)$, that are smooth functions of time. An interesting proposition of this method is that the controls obtained fullfill a necessary (albeit not sufficient) condition for a time-optimal control.
@@
Figure 4: Comparison of GRAPE and MAGICARP execution times to implement 1000 random Haar unitaries. (a) For a $\text{Tb}\text{Pc}_{2}$ molecule. (b) For a $\text{Tb}_{2}\text{Pc}_{3}$ molecule.
@@
It appears numerically that this parametrization yields shorter gate times, all the more as the dimension of the Hilbert space increases. For the triple decker in fig. 4b the smallest execution time found by the MAGICARP is 37 % faster than for the GRAPE method for instance. An important point is that it is increasingly computationally difficult to optimize the $M$ matrix of eq. 39 with the growing size of the Hilbert space. In appendix C, we discuss an optimizer based on the natural gradient which provides solutions below the infidelity threshold for qudits up to $d=16$ dimensions, e.g. the triple decker discussed in Section Y. We used it to obtain the results of fig. 4b. A Julia and a Python code are available on the Github repositories Magicarp and pyMagicarp. Overall, the higher the dimension of the system is, the better the results of the MAGICARP in comparison to GRAPE when comparing median execution times for random gates or the minimum execution time for a specific gate (the QFT) as shown in fig. 5.
@@
Figure 5: Comparison of GRAPE and MAGICARP execution times on linear graphs with 3 to 9 nodes, as a function of the Hilbert space dimension. There is at least 100 initializations for each dimension. We used the natural gradient descent described in appendix C for the green curves, and for systems of dimension below 6, we also used the more computationally expensive L-BFGS-B optimizer from Scipy [35] described by the yellow curve. The ratios indicate how much faster MAGICARP is compared to GRAPE for a given dimension. (a) Ratio of the median execution times for random Haar gates using GRAPE with random initializations over the median time using MAGICARP. (b) Ratio of the minimum execution times for a QFT gate using GRAPE with random initializations over the minimum time using MAGICARP.
@@
We choose to use a QFT gate to compare the minimum execution time for the MAGICARP and the GRAPE method since this gate is not trivial, is key to many important quantum algorithms as explained in section III.1.2, and has execution times similar to a random gate (gaussian-like) as one can see when comparing fig. 4b and fig. 6a. Note that for small dimensions (3 and 4), their is no improvement of the execution times with the natural gradient optimizer used here since the ratios are below 1 as shown in fig. 5 for the corresponding curve. Though if we compare the MAGICARP with the same gradient method as GRAPE (A L-BFGS-B gradient from SciPy [9, 35].), with the same parameters, and using automatic differentiation, we can compute the results at low dimensions (up to 5) and do witness that the MAGICARP is at least faster than GRAPE both in median time for random gates and in minimum time for the QFT gate for this naive implementation. This observation illustrates that the advantage of MAGICARP over GRAPE discussed in this paper leaves room for further improvements of the execution time by carefully designing the algorithm used to optimize the parameter of MAGICARP.
Comparing these methods only by implementing target gates chosen uniformly at random in the unitary group provides insights into some kind of improvement of the execution time “on average”. However, in practice, not all unitary matrices are of interest for the existing quantum algorithms. We therefore compare the GRAPE and MAGICARP methods on the triple decker ($d=16$) by implementing four particular commonly used quantum gates: QFT, T, X and SUMX [37]. The latter generalizes the CNOT gate to qudits. We obtained significantly better results in fig. 6 for the MAGICARP, regarding both median and minimum execution times.
@@
Figure 6: Comparison of GRAPE and MAGICARP execution times for 1000 random initializations on a $\text{Tb}_{2}\text{Pc}_{3}$ molecule. (a) QFT gate. (b) T gate. (c) X gate. (d) SUMX gate.